Reports 1-1 of 1 Clear search Modify search
PEM (Center)
fabian.arellano - 11:32 Wednesday 18 March 2020 (13661) Print this report
Comment to BS frame hammering test (13603)

I took a loot to the event upon Yokozawa-san's request.

In case you don't see the plot below please click here: 13659.

Date and time of plot below: Sunday 2020-03-15 from 4:30:00 to 5:30:00 UTC (from 13:30 to 14:30 JST).

Besides showing the data that confirms the IP tripped my only comment is that I don't remember the other frames to be as delicate as Yokozawa-san's description of the BS' one. This includes VIS postintallation work and PEM work, e.g. installing accelerometers in SR3 upper oplev platform, removing a heavy leaning flange from SRM frame, etc.

I need to pay attention to this situation after the observation campaign and before we open the chamber for VIS work. I wonder what this may sugget about the assembly of the frame.

The plots show the IP actuation reaches  saturation at -20,000 counts in the three coils, which is the limit set in the medm screen. Then oplev signal also goes out of range (beyond ±100 normalized displacement units).

Images attached to this comment
Search Help